Divide: 15/2 ÷ 10
a. 3/4 b. 4/3 c. 75 d. 1/75 e. 5/2

Answer by jessica43(140)

The trick to this problem is to first re-write it and then multiply by the reciprocal.
(15/2)÷ 10=
(15/2)/10=
(15/2)/(10/1)=
These are equal to each other- just written differently.
Now dividing (15/2) by (10/1) is the same as multiplying (15/2) by (1/10). In other words, you want to turn the second fraction (the one you want to divide by) upside down (this is called the reciprocal). 1/10 is the reciprocal of 10/1. Then you multiply the first fraction by the reciprocal.
(15/2)/(10/1)=
(15/2)*(1/10)=
(15*1)/(2*10)=
(15/20)=
3/4 (I simplified by dividing the numerator and denominator by 5)
Your answer is a) 3/4.
You will follow these same steps for other fraction problems.